# Tree-Wasserstein Distance for High Dimensional Data with a Latent Feature Hierarchy

### v2

@@ 4.2 Hyperbolic Diffusion Embedding
In our approach, we first construct a hyperbolic representation following Lin et al. [2023]. We briefly describe the steps of this construction below. The feature diffusion operator $\mathbf{P}\in\mathbb{R}^{m\times m}$ is built according to Eq. (2), where $d(j,j^{\prime})$ is an initial distance between the features $j$ and $j^{\prime}$ embedded in the ambient space $\mathbb{R}^{n}$. In Sec. 5, we compute the initial distance using the cosine similarity in the ambient space, i.e., $d(j,j^{\prime})=1-\frac{\mathbf{X}_{:,j}\cdot\mathbf{X}_{:,j^{\prime}}}{\left\lVert\mathbf{X}_{:,j}\right\rVert_{2}\left\lVert\mathbf{X}_{:,j^{\prime}}\right\rVert_{2}}$, where $\cdot$ is the dot product. We refer to App. D for further discussion on the initial distance metric, the scale parameter $\epsilon$, and the kernel type. Then, we construct the multi-scale diffusion densities $\bm{\phi}^{k}_{j}=\mathbf{P}^{2^{-k}}\mathbf{e}_{j}\in\mathbb{R}^{m}$ for each feature $j\in[m]$, considering a series of diffusion time steps on a dyadic grid $t=2^{-k}$ for $k\in\mathbb{Z}_{0}^{+}$. These densities give rise to a multi-scale view of the features on the dyadic grid. We then embed each feature $j$ at the $k$-th scale by $(j,k)\mapsto\mathbf{z}_{j}^{k}=[(\bm{\psi}_{j}^{k})^{\top},2^{k/2-2}]^{\top}\in\mathbb{H}^{m+1}$ based on the exponential growth of the Poincaré half-space, where $\bm{\psi}^{k}_{j}=\displaystyle\textstyle\sqrt{\bm{\phi}^{k}_{j}}\in\mathbb{R}^{m}$ is the element-wise square root of the diffusion density $\bm{\phi}^{k}_{j}$. The multi-scale embedding of the corresponding diffusion operators in hyperbolic space [Lin et al., 2023] results in distances that are exponentially scaled. This scaling aligns with the structure of a tree distance, effectively establishing a natural hierarchical relation between different diffusion timescales. While it is well-established that diffusion geometry effectively recovers the underlying manifold [Coifman and Lafon, 2006], recent research by Lin et al. [2023] builds on this by showing that propagated densities, when applied with diffusion times on dyadic grids, can reveal hidden hierarchical metric. This method effectively captures local information from exponentially expanding neighborhoods around each point. We note that other hyperbolic embeddings (e.g., Chamberlain et al. [2017], Nickel and Kiela [2017]) could potentially be considered. However, these methods usually assume a (partially) known graph $T$ instead of high-dimensional observational data $\mathbf{X}$, and therefore, we do not opt to use them (see App. E.3.4 for more details).
@@ 4.3 Hyperbolic Diffusion LCA
The hyperbolic LCA of any two points $\mathbf{z}_{j}^{k},\mathbf{z}_{j^{\prime}}^{k}\in\mathbb{H}^{m+1}$ at the $k$-th scale is defined by their Fréchet mean [Fréchet, 1948] as $\mathbf{z}_{j}^{k}\vee\mathbf{z}_{j^{\prime}}^{k}\coloneqq\underset{\mathbf{z}\in\mathbb{H}^{m+1}}{\arg\min}\left(d_{\mathbb{H}^{m+1}}^{2}(\mathbf{z},\mathbf{z}_{j}^{k})+d_{\mathbb{H}^{m+1}}^{2}(\mathbf{z},\mathbf{z}_{j^{\prime}}^{k})\right)$.
@@
The hyperbolic LCA $\mathbf{z}_{j}^{k}\vee\mathbf{z}_{j^{\prime}}^{k}$ in Def. 4.1 has a closed-form solution, given by
$\mathbf{z}_{j}^{k}\vee\mathbf{z}_{j^{\prime}}^{k}=\left[\frac{1}{2}\left(\bm{\psi}_{j}^{k}+\bm{\psi}_{j^{\prime}}^{k}\right)^{\top},\mathtt{proj}(\mathbf{z}_{j}^{k}\vee\mathbf{z}_{j^{\prime}}^{k})\right]^{\top},$
where $\mathtt{proj}(\mathbf{z}_{j}^{k}\vee\mathbf{z}_{j^{\prime}}^{k})=\left\lVert\left[\frac{1}{2}\left(\bm{\psi}_{j}^{k}-\bm{\psi}_{j^{\prime}}^{k}\right)^{\top},2^{k/2-2}\right]^{\top}\right\rVert_{2}$.
@@
Incorporating the multi-scale of the hyperbolic LCAs arranged on a dyadic grid via the HD-LCA facilitates the recovery of the underlying tree relations. A similar approach was presented in Lin et al. [2023], where it was shown that the distance between the embedding $j\mapsto[(\mathbf{z}_{j}^{0})^{\top},\ldots,(\mathbf{z}_{j}^{K_{c}})^{\top}]^{\top}\in\mathcal{M}=\mathbb{H}^{m+1}\times\ldots\times\mathbb{H}^{m+1}$ in a product manifold of Poincaré half-spaces, given by $d_{\mathcal{M}}(j,j^{\prime})=\sum_{k=0}^{K_{c}}~{}2\sinh^{-1}(2^{-k/2+1}\left\lVert\mathbf{z}^{k}_{j}-\mathbf{z}^{k}_{j^{\prime}}\right\rVert_{2})$, recovers $d_{T}(j,j^{\prime})$. However, [Lin et al., 2023] considers a latent hierarchical structure of the samples. It is fundamentally different than the problem we consider in Sec. 4.1, as it does not learn the latent hierarchy of the features explicitly by constructing a tree, and then incorporate it in a meaningful distance between the samples as we do here (see App. F.4 for more details). As shown in App. F.4, this distance is less effective for document and cell classifications.
@@
  function $\mathtt{TWD}\_\mathtt{latent}\_\mathtt{feature}(\mathbf{X},K_{c})$
@@ 4.4 Binary Tree Construction
Equipped with the definitions and properties above, we now present the proposed tree construction in hyperbolic spaces. As any tree can be transformed into a binary tree [Bowditch, 2007, Cormen et al., 2022], we suggest to construct a rooted binary tree $B$ through an iterative merging process. We begin by assigning the features to the leaf nodes because it was shown that the leaves alone contain adequate information to fully reconstruct the tree [Sarkar, 2011] (see App. F.2 for more details). This is not only efficient but also aligns with established TWD [Indyk and Thaper, 2003, Backurs et al., 2020, Le et al., 2019]. Then, pairs of features are merged iteratively by their HD-LCA (Def. 4.3). Specifically, the pair of features that exhibits the highest similarity in the HD-LCA (i.e., the smallest value of $a_{j,j^{\prime}}$ in Prop. 4.3) is merged. Note that only one of the HD-LCA relations $[j_{2}\vee j_{3}\sim j_{1}]_{\mathcal{M}}$, $[j_{1}\vee j_{3}\sim j_{2}]_{\mathcal{M}}$, or $[j_{1}\vee j_{2}\sim j_{3}]_{\mathcal{M}}$ holds in $B$. Once the nodes are merged, the weight of the edge is assigned with the induced $\ell_{1}$ distance between the respective embeddings. This iterative process constructs the binary tree $B$ in a bottom-up manner, similar to single linkage clustering [Gower and Ross, 1969], with the measure of similarity determined by their HD-LCAs and edge weights assigned with the geodesic distance. We summarize the proposed HD binary tree decoding in Alg. 1.
We note that the distance in the embedding [Lin et al., 2023] could be used in existing graph-based algorithms [Alon et al., 1995] or linkage methods [Murtagh and Contreras, 2012] instead of the HD-LCA in the binary tree decoding (Alg. 1). However, these generic algorithms, while providing upper bounds on the distortion of the constructed tree distances, do not build a geometrically meaningful tree when given a tree metric [Sonthalia and Gilbert, 2020], because they do not exploit the hyperbolic geometry of the embedded space. Indeed, as empirically shown in App. E, constructing a feature tree using existing TWDs, graph-based, or linkage approaches with embedding distance is less effective. This may suggest that the information encoded in the multi-scale hyperbolic LCAs is critical for revealing the latent feature tree. In addition, we argue that having a closed-form solution to the basic element of the tree construction HD-LCA in Prop. 4.3 is an important advantage because it directly enables us to present an effective and interpretable algorithm (Alg. 1) for recovering the underlying tree relationships from the learned hyperbolic embedding.
@@
Let $r$ be the root node of $B$ obtained in Alg. 1 and $\mathbf{z}_{r}$ be the corresponding embedding in $\mathcal{M}$. The HD-LCA depth is equivalent to the Gromov product [Gromov, 1987], i.e., $d_{\mathcal{M}}(\mathbf{z}_{r},\overline{\mathbf{h}}_{j,j^{\prime}})\simeq\langle j,j^{\prime}\rangle_{r}$, where $\langle j,j^{\prime}\rangle_{r}=\frac{1}{2}(d_{B}(j,r)+d_{B}(j^{\prime},r)-d_{B}(j,j^{\prime}))$.
@@
The proof of Thm. 4.1 is in App. C. The validity of Thm. 4.1 relies on the assumption that the local Gaussian affinities between features allow us to reveal the hidden hierarchy. Formally, in the proof of Thm. 4.1, we assume that the discrete diffusion operator $\mathbf{P}$ constructed based on the local Gaussian affinities between features in the ambient space $\mathbb{R}^{n}$, in the limit of $m\rightarrow\infty$ and $\epsilon\rightarrow 0$, converges pointwise to the heat kernel of the underlying hierarchical metric space $(\mathcal{H},d_{\mathcal{H}})$. While $T$ is hidden, the diffusion operator $\mathbf{P}$ is accessible and used to construct another tree, $B$, such that, the tree distances of the inferred tree $d_{B}$ and the hidden tree $d_{T}$ are equivalent. This result is stated in Thm. 4.1. In addition, our HD-LCA is a continuous LCA in the product manifold of $(m+1)$-dimensional Poincaré half-spaces that recovers hidden tree relations. It is closely related to the hyperbolic LCA in Chami et al. [2020], where a continuous LCA in a single 2D Poincaré disk was proposed for hierarchical clustering, and a tree was constructed by minimizing the Dasgupta’s cost [Dasgupta, 2016]. However, their method does not recover the latent feature hierarchy, as shown in Sec. 5.
@@ 4.5 Tree-Wasserstein Distance for High-Dimensional Data with a Latent Feature Hierarchy
The proof of Thm. 4.2 is in App. C. Thm. 4.2 shows that our TWD in Eq. (3) recovers the TWD associated with the latent feature tree $T$ without access to $T$. This could be viewed as an unsupervised Wasserstein metric learning between data samples with the latent feature hierarchy. In addition, our intermediate result in Thm. 4.1 is noteworthy, as it represents a form of unsupervised ground metric learning [Cuturi and Avis, 2014] derived from the latent tree underlying the features.
@@
Seemingly, the need to build a distance matrix and apply the eigendecomposition to the resulting kernel limits the applicability to a small number of features. However, using [Shen and Wu, 2022], the proposed TWD can be computed in $O(m^{1.2})$ compared to the naïve implementation that requires $O(mn^{3}+m^{3}\log m)$ (see App. D). Indeed, in Sec. 5, we show applications with thousands of features.
One limitation of our TWD in Eq. (3) is the restriction to samples from the probability simplex. This requirement, which stems from the definitions of OT framework [Monge, 1781, Kantorovich, 1942] and the Wasserstein and TWDs [Villani, 2009, Cuturi, 2013, Indyk and Thaper, 2003, Backurs et al., 2020, Le et al., 2019, Takezawa et al., 2021, Yamada et al., 2022], restricts its applicability. However, a large body of work [Villani, 2009, Cuturi, 2013, Courty et al., 2017, Solomon et al., 2015, Yurochkin et al., 2019, Alvarez-Melis and Jaakkola, 2018] illustrates that this setting is relevant in a broad range of applications. In addition, this limitation does not affect the tree decoding in Alg. 1, as it can recover the latent feature hierarchy from $\mathbb{R}^{n}$, which is in line with existing tree construction methods [Alon et al., 1995, Murtagh and Contreras, 2012]. Note that the problem of finding distances between samples with feature hierarchy and possibly with negative entries has been addressed in Gavish et al. [2010], Ankenman [2014], Mishne et al. [2016, 2017] using different approaches.
A notable property of our TWD in Eq. (3) is that it can be interpreted using the tree-sliced Wasserstein distance (TSWD) [Le et al., 2019]. Computing the TSWD consists of constructing multiple random trees of different depths from $\mathbf{X}$. Then, the TSWD is given by the average of the TWDs corresponding to these trees. Here, considering only a single scale $k$ in Alg. 1 (as detailed in App. F.1) yields the decoded tree $B^{k}$. Similarly to the TSWD, we show that our TWD can be recast as the sum of the TWDs corresponding to $B^{k}$ at multiple scales $k=0,\ldots,K_{c}$.
@@
The proof of Prop. 4.5 is in App. C. Prop. 4.5 implies that our TWD captures the (tree-sliced) information in the multi-scale trees $B^{k}$ efficiently as it circumvents their explicit construction. In addition, we show in Sec. 5 that the our TWD outperforms the TSWD [Le et al., 2019].
We conclude with a few remarks. First, it is important to note that our primary objective is not to recover the exact hidden $T$, but to effectively approximate the TWD $\mathtt{TW}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}},T)$ from the data matrix $\mathbf{X}$ (i.e., the node attributes from the hidden $T$), which we achieve by the proposed TWD $\mathtt{TW}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}},B)$ in Eq. (3) as stated in Thm. 4.2. Second, our TWD can also be used to approximate the OT distance in linear time. For example, suppose we are interested in computing the OT distance based on a given ground tree metric between the features. While the complexity of this computation is typically $O(m^{3}\log m)$ [Pele and Werman, 2009, Bonneel et al., 2011], we can approximate it by the proposed TWD in linear time, as we demonstrate in the runtime analysis in App. E. Finally, our TWD is differentiable, as detailed in App. F.3, in contrast to existing TWD methods. This differentiability equips the proposed TWD with a continuous gradient, significantly enhancing its applicability in gradient-based optimization, a feature that is not typically found in TWD baselines.

### v3

@@ 4.2 Hyperbolic Diffusion Embedding
In our approach, we first construct a hyperbolic representation following Lin et al. (2023). We briefly describe the steps of this construction below. The feature diffusion operator $\mathbf{P}\in\mathbb{R}^{m\times m}$ is built according to Eq. (2), where $d(j,j^{\prime})$ is an initial distance between the features $j$ and $j^{\prime}$ embedded in the ambient space $\mathbb{R}^{n}$. In Sec. 5, we compute the initial distance using the cosine similarity in the ambient space, i.e., $d(j,j^{\prime})=1-\frac{\mathbf{X}_{:,j}\cdot\mathbf{X}_{:,j^{\prime}}}{\|\mathbf{X}_{:,j}\|_{2}\|\mathbf{X}_{:,j^{\prime}}\|_{2}}$, where $\cdot$ is the dot product. We refer to App. D for further discussion on the initial distance metric, the scale parameter $\epsilon$, and the kernel type. Then, we construct the multi-scale diffusion densities $\bm{\phi}^{k}_{j}=\mathbf{P}^{2^{-k}}\mathbf{e}_{j}\in\mathbb{R}^{m}$ for each feature $j\in[m]$, considering a series of diffusion time steps on a dyadic grid $t=2^{-k}$ for $k\in\mathbb{Z}_{0}^{+}$. These densities give rise to a multi-scale view of the features on the dyadic grid. We then embed each feature $j$ at the $k$-th scale by $(j,k)\mapsto\mathbf{z}_{j}^{k}=[(\bm{\psi}_{j}^{k})^{\top},2^{k/2-2}]^{\top}\in\mathbb{H}^{m+1}$ based on the exponential growth of the Poincaré half-space, where $\bm{\psi}^{k}_{j}=\displaystyle\textstyle\sqrt{\bm{\phi}^{k}_{j}}\in\mathbb{R}^{m}$ is the element-wise square root of the diffusion density $\bm{\phi}^{k}_{j}$. The multi-scale embedding of the corresponding diffusion operators in hyperbolic space (Lin et al., 2023) results in distances that are exponentially scaled. This scaling aligns with the structure of a tree distance, effectively establishing a natural hierarchical relation between different diffusion timescales. While it is well-established that diffusion geometry effectively recovers the underlying manifold (Coifman & Lafon, 2006), recent research by Lin et al. (2023) builds on this by showing that propagated densities, when applied with diffusion times on dyadic grids, can reveal hidden hierarchical metric. This method effectively captures local information from exponentially expanding neighborhoods around each point. We note that other hyperbolic embeddings (e.g., Chamberlain et al. (2017); Nickel & Kiela (2017)) could potentially be considered. However, these methods usually assume a (partially) known graph $T$ instead of high-dimensional observational data $\mathbf{X}$, and therefore, we do not opt to use them (see App. E.3.4 for more details).
@@ 4.3 Hyperbolic Diffusion LCA
The hyperbolic LCA of any two points $\mathbf{z}_{j}^{k},\mathbf{z}_{j^{\prime}}^{k}\in\mathbb{H}^{m+1}$ at the $k$-th scale is defined by their Fréchet mean (Fréchet, 1948) as $\mathbf{z}_{j}^{k}\vee\mathbf{z}_{j^{\prime}}^{k}\coloneqq\underset{\mathbf{z}\in\mathbb{H}^{m+1}}{\arg\min}\left(d_{\mathbb{H}^{m+1}}^{2}(\mathbf{z},\mathbf{z}_{j}^{k})+d_{\mathbb{H}^{m+1}}^{2}(\mathbf{z},\mathbf{z}_{j^{\prime}}^{k})\right)$.
@@
The hyperbolic LCA $\mathbf{z}_{j}^{k}\vee\mathbf{z}_{j^{\prime}}^{k}$ in Def. 4.1 has a closed-form solution, given by $\mathbf{z}_{j}^{k}\vee\mathbf{z}_{j^{\prime}}^{k}=\left[\frac{1}{2}\left(\bm{\psi}_{j}^{k}+\bm{\psi}_{j^{\prime}}^{k}\right)^{\top},\mathtt{proj}(\mathbf{z}_{j}^{k}\vee\mathbf{z}_{j^{\prime}}^{k})\right]^{\top}$, where $\mathtt{proj}(\mathbf{z}_{j}^{k}\vee\mathbf{z}_{j^{\prime}}^{k})=\left\lVert\left[\frac{1}{2}\left(\bm{\psi}_{j}^{k}-\bm{\psi}_{j^{\prime}}^{k}\right)^{\top},2^{k/2-2}\right]^{\top}\right\rVert_{2}$.
@@
Incorporating the multi-scale of the hyperbolic LCAs arranged on a dyadic grid via the HD-LCA facilitates the recovery of the underlying tree relations. A similar approach was presented in Lin et al. (2023), where it was shown that the distance between the embedding $j\mapsto[(\mathbf{z}_{j}^{0})^{\top},\ldots,(\mathbf{z}_{j}^{K_{c}})^{\top}]^{\top}\in\mathcal{M}=\mathbb{H}^{m+1}\times\ldots\times\mathbb{H}^{m+1}$ in a product manifold of Poincaré half-spaces, given by $d_{\mathcal{M}}(j,j^{\prime})=\sum_{k=0}^{K_{c}}~{}2\sinh^{-1}(2^{-k/2+1}\left\lVert\mathbf{z}^{k}_{j}-\mathbf{z}^{k}_{j^{\prime}}\right\rVert_{2})$, recovers $d_{T}(j,j^{\prime})$. However, (Lin et al., 2023) considers a latent hierarchical structure of the samples. It is fundamentally different than the problem we consider in Sec. 4.1, as it does not learn the latent hierarchy of the features explicitly by constructing a tree, and then incorporate it in a meaningful distance between the samples as we do here (see App. F.4 for more details). As shown in App. F.4, this distance is less effective for document and cell classifications.
@@
  function $\mathtt{TWD}\_\mathtt{latent}\_\mathtt{HieFeature}(\mathbf{X},K_{c})$
@@ 4.4 Binary Tree Construction
Equipped with the definitions and properties above, we now present the proposed tree construction in hyperbolic spaces. As any tree can be transformed into a binary tree (Bowditch, 2007; Cormen et al., 2022), we suggest to construct a rooted binary tree $B$ through an iterative merging process. We begin by assigning the features to the leaf nodes because it was shown that the leaves alone contain adequate information to fully reconstruct the tree (Sarkar, 2011) (see App. F.2 for more details). This is not only efficient but also aligns with established TWD (Indyk & Thaper, 2003; Backurs et al., 2020; Le et al., 2019). Then, pairs of features are merged iteratively by their HD-LCA (Def. 4.3). Specifically, the pair of features that exhibits the highest similarity in the HD-LCA (i.e., the smallest value of $a_{j,j^{\prime}}$ in Prop. 4.3) is merged. Note that only one of the HD-LCA relations $[j_{2}\vee j_{3}\sim j_{1}]_{\mathcal{M}}$, $[j_{1}\vee j_{3}\sim j_{2}]_{\mathcal{M}}$, or $[j_{1}\vee j_{2}\sim j_{3}]_{\mathcal{M}}$ holds in $B$. Once the nodes are merged, the weight of the edge is assigned with the induced $\ell_{1}$ distance between the respective embeddings. This iterative process constructs the binary tree $B$ in a bottom-up manner, similar to single linkage clustering (Gower & Ross, 1969), with the measure of similarity determined by their HD-LCAs and edge weights assigned with the geodesic distance. We summarize the proposed HD binary tree decoding in Alg. 1.
We note that the distance in the embedding (Lin et al., 2023) could be used in existing graph-based algorithms (Alon et al., 1995) or linkage methods (Murtagh & Contreras, 2012) instead of the HD-LCA in the binary tree decoding (Alg. 1). However, these generic algorithms, while providing upper bounds on the distortion of the constructed tree distances, do not build a geometrically meaningful tree when given a tree metric (Sonthalia & Gilbert, 2020), because they do not exploit the hyperbolic geometry of the embedded space. Indeed, as empirically shown in App. E, constructing a feature tree using existing TWDs, graph-based, or linkage approaches with embedding distance is less effective. This may suggest that the information encoded in the multi-scale hyperbolic LCAs is critical for revealing the latent feature tree. In addition, we argue that having a closed-form solution to the basic element of the tree construction HD-LCA in Prop. 4.3 is an important advantage because it directly enables us to present an effective and interpretable algorithm (Alg. 1) for recovering the underlying tree relationships from the learned hyperbolic embedding.
@@
Let $r$ be the root node of $B$ obtained in Alg. 1 and $\mathbf{z}_{r}$ be the corresponding embedding in $\mathcal{M}$. The HD-LCA depth is equivalent to the Gromov product (Gromov, 1987), i.e., $d_{\mathcal{M}}(\mathbf{z}_{r},\overline{\mathbf{h}}_{j,j^{\prime}})\simeq\langle j,j^{\prime}\rangle_{r}$, where $\langle j,j^{\prime}\rangle_{r}=\frac{1}{2}(d_{B}(j,r)+d_{B}(j^{\prime},r)-d_{B}(j,j^{\prime}))$.
@@
The proof of Thm. 4.1 is in App. C. The validity of Thm. 4.1 relies on the assumption that the local Gaussian affinities between features allow us to reveal the hidden hierarchy. Formally, in the proof of Thm. 4.1, we assume that the discrete diffusion operator $\mathbf{P}$ constructed based on the local Gaussian affinities between features in the ambient space $\mathbb{R}^{n}$, in the limit of $m\rightarrow\infty$ and $\epsilon\rightarrow 0$, converges pointwise to the heat kernel of the underlying hierarchical metric space $(\mathcal{H},d_{\mathcal{H}})$. While $T$ is hidden, the diffusion operator $\mathbf{P}$ is accessible and used to construct another tree, $B$, such that, the tree distances of the inferred tree $d_{B}$ and the hidden tree $d_{T}$ are equivalent. This result is stated in Thm. 4.1. In addition, our HD-LCA is a continuous LCA in the product manifold of $(m+1)$-dimensional Poincaré half-spaces that recovers hidden tree relations. It is closely related to the hyperbolic LCA in Chami et al. (2020), where a continuous LCA in a single 2D Poincaré disk was proposed for hierarchical clustering, and a tree was constructed by minimizing the Dasgupta’s cost (Dasgupta, 2016). However, their method does not recover the latent feature hierarchy, as shown in Sec. 5.
@@ 4.5 Tree-Wasserstein Distance for High-Dimensional Data with a Latent Feature Hierarchy
The proof of Thm. 4.2 is in App. C. Thm. 4.2 shows that our TWD in Eq. (3) recovers the TWD associated with the latent feature tree $T$ without access to $T$. This could be viewed as an unsupervised Wasserstein metric learning between data samples with the latent feature hierarchy. In addition, our intermediate result in Thm. 4.1 is noteworthy, as it represents a form of unsupervised ground metric learning (Cuturi & Avis, 2014) derived from the latent tree underlying the features.
@@
Seemingly, the need to build a distance matrix and apply the eigendecomposition to the resulting kernel limits the applicability to a small number of features. However, using (Shen & Wu, 2022), the proposed TWD can be computed in $O(m^{1.2})$ compared to the naïve implementation that requires $O(mn^{3}+m^{3}\log m)$ (see App. D). Indeed, in Sec. 5, we show applications with thousands of features.
One limitation of our TWD in Eq. (3) is the restriction to samples from the probability simplex. This requirement, which stems from the definitions of OT framework (Monge, 1781; Kantorovich, 1942) and the Wasserstein and TWDs (Villani, 2009; Cuturi, 2013; Indyk & Thaper, 2003; Backurs et al., 2020; Le et al., 2019; Takezawa et al., 2021; Yamada et al., 2022), restricts its applicability. However, a large body of work (Villani, 2009; Cuturi, 2013; Courty et al., 2017; Solomon et al., 2015; Yurochkin et al., 2019; Alvarez-Melis & Jaakkola, 2018) illustrates that this setting is relevant in a broad range of applications. In addition, this limitation does not affect the tree decoding in Alg. 1, as it can recover the latent feature hierarchy from $\mathbb{R}^{n}$, which is in line with existing tree construction methods (Alon et al., 1995; Murtagh & Contreras, 2012). Note that the problem of finding distances between samples with feature hierarchy and possibly with negative entries has been addressed in Gavish et al. (2010); Ankenman (2014); Mishne et al. (2016; 2017) using different approaches.
A notable property of our TWD in Eq. (3) is that it can be interpreted using the tree-sliced Wasserstein distance (TSWD) (Le et al., 2019). Computing the TSWD consists of constructing multiple random trees of different depths from $\mathbf{X}$. Then, the TSWD is given by the average of the TWDs corresponding to these trees. Here, considering only a single scale $k$ in Alg. 1 (as detailed in App. F.1) yields the decoded tree $B^{k}$. Similarly to the TSWD, we show that our TWD can be recast as the sum of the TWDs corresponding to $B^{k}$ at multiple scales $k=0,\ldots,K_{c}$.
@@
The proof of Prop. 4.5 is in App. C. Prop. 4.5 implies that our TWD captures the (tree-sliced) information in the multi-scale trees $B^{k}$ efficiently as it circumvents their explicit construction.
We conclude with a few remarks. First, it is important to note that our primary objective is not to recover the exact hidden $T$, but to effectively approximate the TWD $\mathtt{TW}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}},T)$ from the data matrix $\mathbf{X}$ (i.e., the node attributes from the hidden $T$), which we achieve by the proposed TWD $\mathtt{TW}(\mathbf{x}_{i},\mathbf{x}_{i^{\prime}},B)$ in Eq. (3) as stated in Thm. 4.2. Second, our TWD can also be used to approximate the OT distance in linear time. For example, suppose we are interested in computing the OT distance based on a given ground tree metric between the features. While the complexity of this computation is typically $O(m^{3}\log m)$ (Pele & Werman, 2009; Bonneel et al., 2011), we can approximate it by the proposed TWD in linear time, as we demonstrate in the runtime analysis in App. E. Third, our method applies to general applications with or without latent hierarchical structure (see App. E.4). Finally, our TWD is differentiable, as detailed in App. F.3, in contrast to existing TWD methods. This differentiability equips the proposed TWD with a continuous gradient, significantly enhancing its applicability in gradient-based optimization, a feature that is not typically found in TWD baselines.